2. x+4y= 14 2x - y=1 x=2, y=3 3. 5x + 3y = 1 3x + 4y = -6 x=2, y=-3 | 4, 2y- 6x =7 3x - y=9 No solution/Parallel lines
Answer #1

Solcetion page No ② x+4y= 14 221 by - 1 we conte the system in the foron AX = B where A= [24] , 6={14] 54][3]-[*] form Now Co.by we get, - now by YSR & SRL fo:](%)=[%] we get y = -3 x + 2 y = 33 x + Z (-3) = - 3 - x=28 4 = -3 is sometion for ginen syfrom there weget Lрела е моя oox toy = 252 x- Yo Y = -7/6 but o.xto y = 0 - 2522 Whalf which is contradiction the giuen syste